MICROWAVE INVESTIGATIONS OF C_{5}H_{5}N-SO_{3} AND HCCCN-SO_{3}: THE PRINCIPLE OF HARD AND SOFT ACIDS AND BASES APPLIED TO PARTIALLY BONDED SYSTEMS

\underline{S. W. HUNT}, D. L. FIACCO, M. CRADDOCK, and K. R. LEOPOLD , Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

The Lewis acid-base adducts C_{5}H_{5}N-SO_{3} and HCCCN-SO_{3} have been studied by Fourier transform microwave spectroscopy. The spectrum of C_{5}H_{5}N-SO_{3} indicates a short N-S bond length of 1.91540(66) Å and free rotation of the SO_{3} unit. In contrast, HCCCN-SO_{3} is more weakly bound with an N-S distance of 2.5676(76) Å, only slightly less than the expected van der Waals interaction distance. The NSO angles are 98.9212(45)^{o} and 91.89(36)^{o} for C_{5}H_{5}N-SO_{3} and HCCCN-SO_{3} respectively. A Townes and Dailey analysis of the ^{14}N quadrupole coupling constant of C_{5}H_{5}N-SO_{3} indicates a transfer of 0.54 electrons upon formation of the dative bond. This is a physical measurement of the ”soft” portion of the chemical interaction and comparisons are made with other adducts of SO_{3}. Bonding is considered in light of Pearson’s concept of Hard and Soft Acids and Bases, ^{a} noting the correlation of adduct properties, such as electron transfer and bond length, with the energy gap between the donor and acceptor orbital.

aR. G. Pearson J. Am. Chem. Soc. 85, 3533 (1963).